# Dimensionality reduction techniques to support

### v1

@@ Abstract
Keywords: Dimensionality reduction, Principal component analysis, Autoencoder, Insider trading, Market abuse, Unsupervised learning.
1 Introduction
@@
Insider trading is the unlawful practice of trading by exploiting nonpublic confidential information about a listed company. It is a type of market abuse: it prevents full and effective market integrity, it violates natural demand-supply dynamics, it compromises public confidence. Knowing in advance how the price will likely move in response to the release of confidential information to the market, i.e. price sensitive event (PSE), such as, for example, the announcement of a takeover bid, can be easily exploited to make a profit. Such a type of practice is prohibited or criminalized in most jurisdictions around the world [4]. However, rules are specific of each country and efforts in persecuting insider trading vary considerably. In the European Union, it is expressly prohibited and administratively sanctioned. The member states are left with the possibility of also imposing criminal sanctions.
@@
The “proof” and the subsequent imposition of a sanction (either administrative or criminal) to an investor that has operated as an insider is however a complex process, involving many steps: (i) the detection of alerts pointing to anomalies that appear attributable to abusive behaviors, (ii) the concrete assessment of the allegedly suspicious conduct with respect to possible rationale that may have supported the strategy under analysis, (iii) the investigation phase aimed at gathering evidence and clues of the abusive conduct, and (iv) the subsequent legal trial to confirm the fact that the unlawful conduct was committed.
In [19], we focus on the first two steps. We propose a methodology, based on unsupervised machine learning techniques, that is capable of providing an indication on whether the trading behavior of an investor or a group of investors is anomalous or not, thus supporting the monitoring and surveillance processes by the competent Authority and the assessment of the conduct. Our previous approach combines two methods, which are both based on well-known techniques, the k-means clustering algorithm [12] and the statistically validated co-occurrence networks [24]. With this new work, we want to provide an extension of the first method of [19]. The latter aims at identifying investors with suspicious behavior related to a price-sensitive event, by means of a dynamic clustering approach. This is done by focusing both on the discontinuities in the trading activity of single investors with respect to their normal activity and to the behavior of their peers, i.e., investors with similar activity. The first step of this method is the characterization of the trading activity of each investor in several time windows. Focusing on a time window, each investor is associated with a point in a three-dimensional space, which corresponds to three trading features (signed turnover, magnitudo, maximum exposure) that are relevant to our insider trading detection task and summarize the activity of the investor in that time window. The choice of three features was motivated by explanability reasons, since we can have a graphical representation and more easily interpret our findings. In this new work, we aim to overcome this choice of the three trading features by employing a dimensionality reduction approach. The idea is that the model should identify the features more relevant to investors’ characterization by itself.
1.1 Literature review
Our work suits into the framework of anomaly detection. This field has been widely explored in the literature, especially in the last years, when its developments have been going at the same pace as machine learning. The applications in the field of financial fraud detection are numerous [25] and, among them, some works are related to the detection of market abuse such as [21, 8, 17, 22]. However, insider trading detection in stock markets is a fairly unexplored topic.
Anomaly detection basically consists in identifying data instances that cannot be associated with normal behavior and that are rare in the data set. The goal of anomaly detection is to define a region of the features’ space where normal observations lie; observations that do not lie in this region are defined as anomalies [5]. Identifying this normality region is not straightforward: the boundary between normal and anomalous behavior is not always sharp, behaviors that are actually anomalous could be disguised in order not to be identified, the definition of normal behavior could be time varying and it is strongly dependent on the application domain, it is difficult to distinguish noise from anomalous behavior [5]. From a practical point of view, there are four main aspects which determine the formulation of the anomaly detection method: availability of data labels, the desired output of the technique, the nature of the input data, the type of anomaly. A different type of anomaly detection approach is employed depending on the availability of data labels: supervised when each observation is labeled as normal or anomalous, semi-supervised when training data do not contain any anomalies and unsupervised when no labels are provided as in our interest case. Typically, the output of the anomaly detection algorithms associates with each observation a score, which quantifies the magnitude of its anomalous character. Setting a suited threshold, the ranked list of anomalies can provide labels for each data instance. Data instances can be of various type (binary/categorical/continuous, univariate/multivariate) and independent among them or related to each other, as it is the case of time series and sequences, spatial data and graph data, for which ad hoc methodologies have to be employed [2, 1]. Concerning the type of anomalies, the standard case is represented by point anomalies, which are single elements identified as anomalous; they could be global or local depending on whether the entire feature space or a specific region of it is considered [10]. Interestingly, there are cases where an element can be seen as normal, but when a given context is taken into account, it turns out to be an anomaly. We refer to this type as contextual anomalies [5], also termed conditional anomalies [23]. It may happen, for instance, that an investor has operated on a stock and, without a context, such an operation looks similar to other operations in the market. However, when compared to the own past behavior of the investor or to the operations of other investors, some discontinuity or synchronization patterns may be revealed and the operation could turn out to be identified as anomalous. Contextual anomalies problems can be tackled by algorithms for point anomaly detection once the context is included as a new feature. Finally, we could have data instances that are normal if considered individually, while they are anomalous together: they are the so-called collective anomalies [5] and can occur in a data set where data instances are dependent.
@@ 1.1 Literature review
In this work, we develop an approach which aims at identifying contextual anomalies and lies in the reconstruction-based paradigm [20]. This framework aims at training models that reconstruct normal data instances well. In this way, we expect that anomalous data will be reconstructed with a large error. An anomaly is detected when the reconstruction error is greater than a threshold i.e.
@@
The standard model employed in reconstruction-based approaches is Principal Component Analysis (PCA) [14]. Its goal is to obtain a compressed representation of data, retaining the most important features. Data are mapped to a lower dimensional space by orthogonal transformations that aim at maximizing data variance or equivalently, minimizing reconstruction error.
The nonlinear counterpart of PCA is an autoencoder [11]. As PCA, autoencoders’s goal is to minimize reconstruction errors but, in this case, the compression and decompression steps are made by means of neural network layers. Complex autoencoder architectures can be devised, as deep, convolutional, LSTM, variational autoencoders [11]. Also, generative adversarial networks have been used in reconstruction-based methods [3]. Moreover, combined approaches as in [6] can be employed.
@@
The paper is organized as follows. In Section 2, the proposed method is described. Section 3 presents the data set we use in our empirical analysis and Section 4 presents the results obtained by our method, with a special focus on one PSE. Finally, conclusions are drawn in Section 5. In the appendix, some figures, which are explained in the main text, are reported, and other collateral issues are investigated.
2 Method
2.1 Overview
@@
$X\rightarrow Z=f_{1}(X)\rightarrow\hat{X}=f_{2}(Z)\text{ where }X,\hat{X}\in\mathbb{R}^{N,T},\ Z\in\mathbb{R}^{N,K}$
and such that the reconstruction error is minimized with respect to the transformations $f_{1}$ and $f_{2}$ i.e.
@@
  • compute $n_{t}=card\{i:t_{i}^{*}=t\}\ \forall t=1,\ldots,T$, where $n_{t}$ is the number of investors having the largest reconstruction error on day $t$;
@@
Finally, in order to detect potential insiders, we devise a method which is based on the following idea. In order to be anomalous, an investor should satisfy the following conditions: (1) to have at least one day for which the reconstruction error is large, (2) the corresponding time lies in the investigation period, (3) she has either a small number of activity days or her identified anomalous score is on a day when not too many other investors do, and (4) she is in a net buying position on the day of the PSE. This last condition can be set by imposing that the difference between the position on the PSE and the position on the first day of the reference period is larger than a threshold, that we choose equal to $0.5$.
@@
This criterion depends on three threshold parameters $d_{\theta}$, $\epsilon_{\theta}$, and $n_{\theta}$. The parameter $d_{\theta}$, the minimal number of days in item (3) of the criterion above, is set to $3$. Instead, we choose $\epsilon_{\theta}$ and $n_{\theta}$ in a data driven fashion: we estimate the probability density function of the anomaly scores $s^{*}_{i}$ and of the times of the largest reconstruction errors $t^{*}_{i}$. Since we observe that the former distribution is bimodal (see the left panel of Figure 3), we expect that normal investor profiles are associated to small anomaly scores $s_{i}^{*}$, while anomalous ones to high scores. Thus $\epsilon_{\theta}$ is chosen as the local minimum between the two modes of the distribution. In practice, we relied on the module signal of the Python library scipy. Finally, $n_{\theta}$ is chosen as the top decile of its distribution.
@@
2.2 Dimensionality reduction methods
Principal Component Analysis
@@ 2.2 Dimensionality reduction methods
$X\rightarrow Z_{K}=XP_{K}\rightarrow\hat{X}=Z_{K}P_{K}^{T}\text{ where }X,\hat{X}\in\mathbb{R}^{N,T},\ Z_{K}\in\mathbb{R}^{N,K},\ P_{K}\in\mathbb{R}^{T,K},$
and the transformation matrix $P_{K}$ is such that the reconstruction error is minimized with a rank constraint, i.e.
$\hat{X}=\arg\min_{\hat{X}^{\prime}:\ rank(\hat{X}^{\prime})\leq K}||X-\hat{X}^{\prime}||_{F}^{2}.$
@@
$Cov(X)=\frac{1}{N}X^{T}X=P\Lambda P^{T}$
where $P\in\mathbb{R}^{T,T}$ is orthogonal ($P^{T}=P^{-1}$), $\Lambda\in\mathbb{R}^{T,T}$ and $\Lambda=diag(\lambda_{1},\lambda_{2},\ldots,\lambda_{T})$ with $\lambda_{1}>\lambda_{2}>\ldots>\lambda_{T}$. The sum of the eigenvalues $\lambda_{1},\ldots,\lambda_{T}$ of the covariance matrix is the total variance of the data. Thus, keeping more components means being able to explain more data variability. The eigenvectors matrix $P$ is defined as $P=[p_{1},p_{2},\ldots,p_{T}]$ and $p_{i}$, $i=1,\ldots,T$ are the loading vectors or principal components. The dimensionality reduction with $K$ components is obtained as
@@
It is evident that PCA is a decorrelation transformation and that its solution is not unique. Indeed, the loss is invariant under the transformation $P\rightarrow PU$, where $U$ is any orthogonal matrix. Under this transformation, the loading vectors are transformed into a different orthonormal basis for the same subspace. Moreover, according to the Eckart-Young theorem [9], the truncated SVD is the best low-rank approximation of a matrix. Therefore, PCA is the linear dimensionality reduction method that minimizes the least squares error in the distortion when we project back to the original space [^9] Observe that our starting data set is in the format $N\times T$, i.e. the trading days are the features that are subjected to compression. Alternatively, we could start with a data set $Y\in\mathbb{R}^{T,N}$ where the features are the investors. In this case, PCA consists in identifying a subset of investors, or combinations of them, with a major role in the characterization of agents’ trading activity. However, this would lead to a more time consuming and computationally expensive procedure. Indeed, we would need to obtain the eigendecomposition of the covariance matrix $Cov(Y)\in\mathbb{R}^{N,N}$, which is much larger than $Cov(X)\in\mathbb{R}^{T,T}$ since $N\gg T$. Furthermore, as shown in Appendix B, if the same feature scaling is applied on the data, the results we obtain starting from the data set in two different formats are analogous..
Finally, it is worth noticing that, compared to other insider detection methods, the one based on PCA does not depend on the knowledge of the initial position of the investors. As explained in Subsection 2.1, this information is indeed lacking in our dataset and in general it can be difficult to obtain because it requires the knowledge of the whole past history of an investor trading activity. Clustering based on k-means, adopted in [19] depends instead on the arbitrary choice of the initial position of the investors. To show that this is not the case for PCA, we prove that the PCA reconstructed position error $\epsilon$ is invariant under the addition of an arbitrary constant to the vector of positions of a given investor. Proof. Let us consider the vector $x_{i}$ describing the position of investor $i$ and let us add an arbitrary constant $C$ to it, obtaining $x_{i}^{C}=x_{i}+C\mathbb{I}_{Tx1}$. Denoting with $\hat{x_{i}}$ and $\hat{x}_{i}^{C}$ the reconstructed positions, it is direct to show that $\hat{x_{i}}^{C}=\hat{x}_{i}+C\mathbb{I}_{Tx1}$. First, we observe that, given the high number of investors, performing PCA on a dataset where investor $i$ has position $x_{i}$ and then, on a dataset where investor $i$ has position $x_{i}^{C}$ leads to loading vectors which are basically the same.
@@
$\begin{split}\hat{x}_{i}^{C}(t)=&\sum_{k=1}^{K}(x_{i}^{C}\cdot p_{k})p_{k}(t)=\sum_{k=1}^{K}[(x_{i}+C\mathbb{I}_{Tx1})\cdot p_{k}]p_{k}(t)=\\ =&\sum_{k=1}^{K}(x_{i}\cdot p_{k})p_{k}(t)+C\sum_{k=1}^{K}(\mathbb{I}_{Tx1}\cdot p_{k})p_{k}(t)=\\ &=\hat{x_{i}}(t)+C\big{\{}[p_{1}(1)p_{1}(t)+p_{2}(1)p_{2}(t)+\ldots+p_{K}(1)p_{K}(t)]+\ldots+\\ &+[p_{1}(t)p_{1}(t)+p_{2}(t)p_{2}(t)+\ldots+p_{K}(t)p_{K}(t)]+\ldots+\\ &+[p_{1}(T)p_{1}(t)+p_{2}(T)p_{2}(t)\ldots+p_{K}(T)p_{K}(t)]\big{\}}=\\ &=\hat{x_{i}}(t)+C[(P_{K}P_{K}^{T})_{1t}+\ldots+(P_{K}P_{K}^{T})_{tt}+\ldots+(P_{K}P_{K}^{T})_{Tt}]=\\ &=\hat{x_{i}}(t)+C\ \ \ \ \ \ \forall t=1,\ldots,T\end{split}$
where $(P_{K}P_{K}^{T})_{lm}$ is the element $l,m$ of the matrix $P_{K}P_{K}^{T}$ and in the last step we exploit that the matrix $P_{K}$ is orthogonal. As explained in Subsection 2.1, our anomaly detection approach is reconstruction-based and $\epsilon_{i}=\epsilon_{i}^{C}$ indeed,
@@
$\begin{split}\hat{x}_{i}^{C}(t)=&\sum_{k=1}^{K}(x_{i}^{C}\cdot p_{k})p_{k}(t)=\sum_{k=1}^{K}\Bigg{(}\frac{\psi_{i}+C\mathbb{I}_{Tx1}}{||\psi_{i}+C\mathbb{I}_{Tx1}||_{\infty}}\cdot p_{k}\Bigg{)}p_{k}(t)=\\ =&\sum_{k=1}^{K}\Bigg{(}\frac{\psi_{i}}{||\psi_{i}+C\mathbb{I}_{Tx1}||_{\infty}}\cdot p_{k}\Bigg{)}p_{k}(t)+C\sum_{k=1}^{K}\Bigg{(}\frac{\mathbb{I}_{Tx1}}{||\psi_{i}+C\mathbb{I}_{Tx1}||_{\infty}}\cdot p_{k}\Bigg{)}p_{k}(t)=\\ =&\frac{||\psi_{i}||_{\infty}}{||\psi_{i}+C\mathbb{I}_{Tx1}||_{\infty}}\sum_{k=1}^{K}\Bigg{(}\frac{\psi_{i}}{||\psi_{i}||_{\infty}}\cdot p_{k}\Bigg{)}p_{k}(t)+\frac{C}{||\psi_{i}+C\mathbb{I}_{Tx1}||_{\infty}}=\\ =&\frac{||\psi_{i}||_{\infty}}{||\psi_{i}+C\mathbb{I}_{Tx1}||_{\infty}}\hat{x}_{i}(t)+\frac{C}{||\psi_{i}+C\mathbb{I}_{Tx1}||_{\infty}}\ \ \ \ \ \ \forall t=1,\ldots,T,\end{split}$
and the reconstruction error is
@@
which holds if $\max_{t}{|\psi_{i}(t)|}=\max_{t}{|\psi_{i}^{C}(t)|}$ given that $\max_{t}|\psi_{i}^{C}(t)|\neq 0$. As we will show in Section 4, this last condition holds for the majority of the profiles in our dataset.
Autoencoders

### v2

@@ Abstract
Machine learning, Time series analysis, Decision support, Principal component analysis, Autoencoder, Insider trading, Market abuse, Unsupervised learning.
I Introduction
@@
Insider trading is the unlawful practice of trading by exploiting nonpublic confidential information about a listed company. It is a type of market abuse: it prevents full and effective market integrity, it violates natural demand-supply dynamics, it compromises public confidence. The announcement of confidential information to the market i.e. price sensitive events (PSEs), implies a specific price response. For instance, after a PSE, as the announcement of a takeover bid, the price is likely to increase. Therefore, foreknowing a PSE can be easily exploited to make a profit.
Most jurisdictions around the world prohibit or criminalize such a type of practice [1]. However, each country has specific rules and enforcement. In the European Union, insider trading is expressly prohibited and administratively sanctioned (Directive 2014/57/EU so-called MAD II, European Regulation 596/2014 so-called MAR) [2]. In Italy the first Market Abuse Directive (MAD) passed in 2005, thereby introducing stringent administrative sanctions in addition to the pre-existing criminal penalties [3].
@@
However, the “proof” and the subsequent imposition of a sanction (either administrative or criminal) to an investor that has operated as an insider is a complex process, involving many steps: (i) the detection of alerts pointing to anomalous behaviors, (ii) the concrete assessment of the allegedly suspicious conduct, (iii) the investigation phase aimed at gathering evidence and clues of the abusive conduct, and (iv) the subsequent legal trial to confirm the unlawful conduct.
As part of an ongoing collaboration between academia and the Italian market supervising authority (Consob), in [4] we recently proposed a methodology which, focusing on the first and partially second step above, supports the monitoring and surveillance processes by the competent Authority and the assessment of the trading conduct. This is done by focusing on the discontinuities in the trading activity of single investors with respect both to their normal activity and to the behavior of their peers, i.e. investors with similar activity. The peers are identified by using an unsupervised dynamic clustering approach based on k-means which takes as input three trading features (termed signed turnover, magnitudo, maximum exposure) obtained from the trading time series of each investor. The choice of the features was motivated by their financial meaning and by explanability reasons, since one can easily interpret which features are responsible for the anomalous behavior.
However, choosing to represent a long time series with three scalar features is somewhat arbitrary and might hide important information useful to identify anomalies in the trading activity of an investor. For this reason, in this paper we present a new approach aiming to overcome this limitation by employing dimensionality reduction methods, namely Principal Component Analysis and Autoencoders. This approach maps the whole time series in a set of coordinates in a suitable latent space which retains a large part of its variability. Large deviations between the original time series and the one reconstructed from the latent space indicate potential anomalies, following the so-called reconstruction-based paradigm. In summary, the idea is that the model should learn by itself the features that are more relevant to investors’ characterization. The methodology is tested on a unique dataset, provided by Consob and containing the trading activity of each investor in the Italian stock market in the proximity of several PSEs.
I-A Related literature
Our work suits into the framework of anomaly detection. This field has been widely explored in the literature, especially in recent years, when its developments have been moving at the same pace as machine learning. The applications in the field of financial fraud detection are numerous [5] and, among them, several works are related to the detection of market abuse such as [14, 16, 18, 9, 12, 13, 6, 15, 17, 7, 8, 10, 11]. In particular, [14, 16, 18, 9, 12, 6, 7, 10, 11] investigate insider trading. By employing supervised approaches, they focus on the identification of market conditions prone to insider trading and on the characterization of insiders’ trading behaviors (for more details, refer to the papers and to [4] for a brief summary of them). Our previous work [4] as well as this new paper, differ from previous literature because of our goal and unavailability of labels. The latter force us to rely on an unsupervised approach. Concerning the goal, we aim to identify investors engaging in suspicious activities related to PSEs. This task is feasible thanks to the transaction reporting database provided to us. This dataset comprehensively covers the daily activity of all investors trading Italian stocks over a period of approximately three years, enabling us to effectively track the trading activity of each individual investor.
Anomaly detection amounts at identifying data instances that deviate from normal behavior and that are uncommon within the data set. Its aim is to identify a region of the features’ space where normal observations lie; observations that do not belong to this region are defined as anomalies [19]. Determining this normality region is not straightforward and is accompanied by several difficulties. The formulation of an anomaly detection method involves four main aspects [19]:
  1. 1.
    the availability of data labels causes the employment of a different approach: supervised when each observation is labeled as normal or anomalous, semi-supervised when training data do not contain any anomalies and unsupervised when no labels are provided as in our interest case;
  2. 2.
    the desired output of the technique is usually a score, which is associated with each observation and quantifies the magnitude of its anomalous character. Setting a suited threshold, each data instance is labelled;
  3. 3.
    input data can be of various type (binary, categorical, continuous, univariate, multivariate) and independent among them or related to each other. Time series and sequences, spatial data and graph data are examples of dependent data and ad hoc methodologies have to be employed for them [21, 20];
  4. 4.
    anomalies can manifest in different forms. Point anomalies are single elements and they could be global or local depending on whether the entire feature space or a specific region of it is considered. Contextual anomalies, also termed conditional anomalies [22] require that a given context is taken into account in order to identify them. Finally, collective anomalies have data instances that are anomalous if they are considered together and they can occur in a data set where data instances are dependent.
@@ I-A Related literature
In this work, we develop an approach which aims at identifying contextual anomalies by relying on continuous and multivariate data instances and lies in the reconstruction-based paradigm of unsupervised anomaly detection [23]. This framework aims at training models that reconstruct normal data instances well. In this way, we expect that anomalous data will be reconstructed with a large error. An anomaly is detected when the reconstruction error is greater than a threshold i.e.
@@
The standard model employed in reconstruction-based approaches is Principal Component Analysis (PCA). Its goal is to obtain a compressed representation of data, retaining the most important features. Data are mapped to a lower dimensional space by orthogonal transformations that aim at maximizing data variance or equivalently, minimizing reconstruction error [24]. Also variants of this method have been employed in the anomaly detection literature e.g. [25].
The nonlinear counterpart of PCA is an autoencoder [26]. As PCA, autoencoders’s goal is to minimize reconstruction errors but, in this case, the compression and decompression steps are made by means of neural network layers. Multiple are the approaches proposed in the anomaly detection literature which rely on autoencoders e.g. [27]. Complex autoencoder architectures can be devised, as deep, convolutional, LSTM, variational autoencoders [28]. Also, generative adversarial networks have been employed [29, 30] as well as combined approaches [31].
@@
The paper is organized as follows. In Section II, the proposed method is described. Section III presents the data set we use in our empirical analysis and Section IV presents the results obtained by our method, with a special focus on one PSE. Finally, conclusions are drawn in Section V. In the Appendices, some figures, which are explained in the main text, are reported, and other collateral issues are investigated.
II Method
II-A Overview
@@
$X\rightarrow Z=f_{1}(X)\rightarrow\hat{X}=f_{2}(Z)$
where $X,\hat{X}\in\mathbb{R}^{N,T},\ Z\in\mathbb{R}^{N,K}$ and such that the reconstruction error is minimized with respect to the transformations $f_{1}$ and $f_{2}$ i.e.
@@
  • compute $n_{t}=\text{card}\{i:t_{i}^{*}=t\}\ \forall t=1,\ldots,T$, where $n_{t}$ is the number of investors having the largest reconstruction error on day $t$;
@@
Finally, in order to detect potential insiders, we devise a method which is based on the following idea. In order to be anomalous, an investor should satisfy the following conditions: (a) to have at least one day for which the reconstruction error is large, (b) the corresponding time lies in the investigation period, (c) she has either a small number of activity days or her identified anomalous score is on a day when not too many other investors do, and (d) she is in a net buying position on the day of the PSE. This last condition can be set by imposing that the difference between the position on the PSE and the position on the first day of the reference period is larger than a threshold, that we choose equal to $0.5$.
@@
This criterion depends on three threshold parameters $d_{\theta}$, $\epsilon_{\theta}$, and $n_{\theta}$. The parameter $d_{\theta}$, the minimal number of days in item (c) of the criterion above, is set to $3$. Instead, we choose $\epsilon_{\theta}$ and $n_{\theta}$ in a data driven fashion: we estimate the probability density function of the anomaly scores $s^{*}_{i}$ and of the times of the largest reconstruction errors $t^{*}_{i}$. Since we observe that the former distribution is bimodal (see the left panel of Figure 3), we expect that normal investor profiles are associated to small anomaly scores $s_{i}^{*}$, while anomalous ones to high scores. Thus $\epsilon_{\theta}$ is chosen as the local minimum between the two modes of the distribution. In practice, we relied on the module signal of the Python library scipy. Finally, $n_{\theta}$ is chosen as the top decile of its distribution.
@@
II-B Dimensionality reduction methods
II-B1 Principal Component Analysis
@@
$X\rightarrow Z_{K}=XP_{K}\rightarrow\hat{X}=Z_{K}P_{K}^{T}$
where $X,\hat{X}\in\mathbb{R}^{N,T},\ Z_{K}\in\mathbb{R}^{N,K},\ P_{K}\in\mathbb{R}^{T,K}$, and the transformation matrix $P_{K}$ is such that the reconstruction error is minimized with a rank constraint, i.e.
$\hat{X}=\arg\min_{\hat{X}^{\prime}:\ \text{rank}(\hat{X}^{\prime})\leq K}||X-\hat{X}^{\prime}||_{F}^{2}.$
@@
$\text{Cov}(X)=\frac{1}{N}X^{T}X=P\Lambda P^{T}$
where $P\in\mathbb{R}^{T,T}$ is orthogonal ($P^{T}=P^{-1}$), $\Lambda\in\mathbb{R}^{T,T}$ and $\Lambda=\text{diag}(\lambda_{1},\lambda_{2},\ldots,\lambda_{T})$ with $\lambda_{1}>\lambda_{2}>\ldots>\lambda_{T}$. The sum of the eigenvalues $\lambda_{1},\ldots,\lambda_{T}$ of the covariance matrix is the total variance of the data. Thus, keeping more components means being able to explain more data variability. The eigenvectors matrix $P$ is defined as $P=[p_{1},p_{2},\ldots,p_{T}]$ and $p_{i}$, $i=1,\ldots,T$ are the loading vectors or principal components. The dimensionality reduction with $K$ components is obtained as
@@
It is evident that PCA is a decorrelation transformation and that its solution is not unique. Indeed, the loss is invariant under the transformation $P\rightarrow PU$, where $U$ is any orthogonal matrix. Under this transformation, the loading vectors are transformed into a different orthonormal basis for the same subspace. Moreover, according to the Eckart-Young theorem [33], the truncated SVD is the best low-rank approximation of a matrix. Therefore, PCA is the linear dimensionality reduction method that minimizes the least squares error in the distortion when we project back to the original space [^2] Observe that our starting data set is in the format $N\times T$, i.e. the trading days are the features that are subjected to compression. Alternatively, we could start with a data set $Y\in\mathbb{R}^{T,N}$ where the features are the investors. In this case, PCA consists in identifying a subset of investors, or combinations of them, with a major role in the characterization of agents’ trading activity. However, this would lead to a more time consuming and computationally expensive procedure. Indeed, we would need to obtain the eigendecomposition of the covariance matrix $\text{Cov}(Y)\in\mathbb{R}^{N,N}$, which is much larger than $\text{Cov}(X)\in\mathbb{R}^{T,T}$ since $N\gg T$. Furthermore, as shown in Appendix B, if the same feature scaling is applied on the data, the results we obtain starting from the data set in two different formats are analogous..
Finally, it is worth noticing that, compared to other insider detection methods, the one based on PCA does not depend on the knowledge of the initial position of the investors. As explained in Subsection II-A, this information is indeed lacking in our dataset and in general it can be difficult to obtain because it requires the knowledge of the whole past history of an investor trading activity. Clustering based on k-means, adopted in [4] depends instead on the arbitrary choice of the initial position of the investors. To show that this is not the case for PCA, we prove that the PCA reconstructed position error $\epsilon$ is invariant under the addition of an arbitrary constant to the vector of positions of a given investor. Proof. Let us consider the vector $x_{i}$ describing the position of investor $i$ and let us add an arbitrary constant $C$ to it, obtaining $x_{i}^{C}=x_{i}+C\mathbb{I}_{Tx1}$. Denoting with $\hat{x_{i}}$ and $\hat{x}_{i}^{C}$ the reconstructed positions, it is direct to show that $\hat{x_{i}}^{C}=\hat{x}_{i}+C\mathbb{I}_{Tx1}$. First, we observe that, given the high number of investors, performing PCA on a dataset where investor $i$ has position $x_{i}$ and then, on a dataset where investor $i$ has position $x_{i}^{C}$ leads to loading vectors which are basically the same.
@@
$\begin{split}\hat{x}_{i}^{C}(t)=&\sum_{k=1}^{K}(x_{i}^{C}\cdot p_{k})p_{k}(t)=\\ =&\sum_{k=1}^{K}(x_{i}\cdot p_{k})p_{k}(t)+C\sum_{k=1}^{K}(\mathbb{I}_{Tx1}\cdot p_{k})p_{k}(t)=\\ &=\hat{x_{i}}(t)+C\sum_{t^{\prime}=1}^{T}(P_{K}P_{K}^{T})_{t^{\prime}t}\\ &=\hat{x_{i}}(t)+C\ \ \ \ \ \ \forall t=1,\ldots,T\end{split}$
where $(P_{K}P_{K}^{T})_{t^{\prime}t}$ is the element $t^{\prime},t$ of the matrix $P_{K}P_{K}^{T}$ and in the last step we exploit that the matrix $P_{K}$ is orthogonal. As explained in Subsection II-A, our anomaly detection approach is reconstruction-based and $\epsilon_{i}=\epsilon_{i}^{C}$ indeed,
@@
$\begin{split}\hat{x}_{i}^{C}(t)=\frac{||\psi_{i}||_{\infty}}{||\psi_{i}+C\mathbb{I}_{Tx1}||_{\infty}}\hat{x}_{i}(t)+\frac{C}{||\psi_{i}+C\mathbb{I}_{Tx1}||_{\infty}}\end{split}$
$\forall t=1,\ldots,T,$ and the reconstruction error is
@@
which holds if $\max_{t}{|\psi_{i}(t)|}=\max_{t}{|\psi_{i}^{C}(t)|}$ given that $\max_{t}|\psi_{i}^{C}(t)|\neq 0$. As we will show in Section IV, this last condition holds for the majority of the profiles in our dataset.
II-B2 Autoencoders